OpenStudy (anonymous):

the speed of a passenger train is 10 mph faster than the speed of a freight train. the passenger train travels 250 miles in the same time it takes the freight train to travel 200 miles. find the speed of each train

OpenStudy (anonymous):

If the sped of the freight train is F, then the speed of the passenger train is F+10. The time for each is the same. Using R*T=D, we can find the time for each train and equate them, eg, \[250/(F+10)=200/F\]You can solve by "cross-multiplying." Do you know how to proceed from here?

OpenStudy (anonymous):

Solving you get\[250F=200F+2000\]\[50F=2000\]\[F=40\]Which is the rate for the freight train. The passenger train's rate is 40+10 or 50 mph.